OpenStudy (anonymous):

jade ran 6 times her neighborhood to complete a total of 1 mile how many times will she need to run to complete 5/6 of a mile

OpenStudy (anonymous):

6/1 = ?/5/6 5/6 *6/1=30/6 30/6=5 5/1= 5 answer= 5